OpenStudy (anonymous):

What is 30% of 20% of 1,000?

OpenStudy (anonymous):

20% of 1000 is 0.20 * 1000 = 200 Now take 30% of 200

OpenStudy (anonymous):

So if 20% of 1000 is 200 and 30% of 1000 is 300, what happens to the rest of the money?

OpenStudy (anonymous):

The remaining 50%? 200 + 300 = 500 = 50% of 1000

OpenStudy (anonymous):

The question asked what is 30% of 20% of 1000. That is equivalent to 0.3 * 0.2 * 1000

OpenStudy (kaederfds):

60

OpenStudy (zzr0ck3r):

.30 * .20 * 1000

OpenStudy (anonymous):

" of " in math problems means multiply. Remember that, for future math problems. Turn your percents to decimals... 30 % = .30 20 % = .20 so your problem now says.... .30 x .20 x 1000 = your answer :)

OpenStudy (anonymous):

\[\frac{30}{100}*\frac{20}{100}*1000=60 \]